### v1

# Stochastic Periodic Solutions for Newtonian Systems via Lyapunov Function

Abstract
This paper establishes an existence theory for distributed periodic solutions to Newton’s equation with stochastic time-periodic forcing, where the friction matrix is the Hessian of a twice continuously differentiable friction function. Employing the Khasminskii criterion and Lyapunov functions, we prove its existence under the assumptions that both the friction and potential functions tend to positive infinity at infinity, and their gradient inner product grows at least like an even-power polynomial. Provided the potential grows sufficiently fast at infinity, this result persists under bounded friction matrix perturbations. This largely confirms the stochastic Levinson conjecture proposed in [9] and resolves the open problem and the interesting problem posed in [9, P. 342].
Distributed periodic solutions exist in two key cases:
  1. 1.
    Polynomial Case: Both functions are polynomials with positive definite leading terms (a minimal condition for dissipation).
  2. 2.
    Plasma Physics Case: The potential force is bounded, while the friction function has sufficiently high growth (as in many plasma physics models).
Keywords: Stochastic time-periodic Newtonian systems; Stochastic Levinson’s conjecture; Damping Hamiltonian systems; Periodic solutions in distribution; Lyapunov’s method
1 Introduction
In physics, it has been a fundamental principle that a Newtonian system
$\ddot{x}+A(x,\dot{x},t)\dot{x}+\nabla V(x)=e(t),$ (1.1)
with friction must have a periodic solution. That is, the system (1.1) admits a periodic solution if the friction matrix $A(x,\dot{x},t)$ is symmetric and positive definite and $A(x,\dot{x},t),e(t)$ are periodic in time $t$. Mathematically in 1940s, Levinson [14] conjectured the principle is true, which brought important dissipativeness theory (see [15, 2, 23, 24] and references therein). For deterministic Newtonian systems, Li et al. [15] verified Levinson’s conjecture by [15, Theorem 4.2, P. 2619]:
Theorem 1.1
Let $V\in C^{2}(\mathbb{R}^{n})$, $e(t)$ and $A(x,\dot{x},t)$ be continuous and periodic in time $t$ with period $\mathcal{T}$ and $A(x,\dot{x},t)$ locally Lipschitz continuous with respect to $(x,\dot{x})$. Suppose that
(i) $A(x,\dot{x},t)$ is symmetric;
(ii) there exist positive constants $\alpha,\eta,\gamma$ and $l$ such that
$A(x,\dot{x},t)\geq\alpha I_{n};\quad x^{\top}\nabla V(x)\geq\eta\|x\|^{2}\,\vee\,\gamma\,x^{\top}A(x,\dot{x},t)\,x,\qquad\hbox{for}~{}\,\|x\|\geq l,$ (1.2)
where $I_{n}$ is the $n\times n$ identity matrix and $\top$ denotes the transposition.
Then system (1.1) admits at least a $\mathcal{T}$-periodic solution.
(1.2) implies that for every fixed $x$, $A(x,\dot{x},t)$ is bounded with respect to $\dot{x}$. That is, the unboundedness of the friction stems from the variable $x$, which warrants greater attention to the situation of unbounded friction. Note that condition (1.2) implies the potential’s growth rate exceeds that of the friction by more than 2. However, this does not hold for the well-known van der Pol equation under time-periodic forcing.
In many contexts, systems are subject to stochastic noise and time-periodic recurrence, which can be modeled as stochastic time-periodic Newtonian systems. From a statistical perspective, Jiang, Li, and Yang [9] utilized the concept of periodicity in distribution — a natural framework for describing periodic phenomenon in stochastic processes. Building on this, they proposed the following stochastic analogue of Levinson’s conjecture.
Conjecture: Stochastic time-periodic Newtonian system
$\ddot{x}+A(x,\dot{x},t)\dot{x}+\nabla V(x)=\dot{B}_{t}+e(t),$ (1.3)
admits $\mathcal{T}$-periodic solution in distribution if this system is dissipative.
Using the Wong-Zakai approximation method, Lyapunov’s method, and Horn’s fixed point theorem, the authors proved this conjecture holds if the potential function $V(x)$ exhibits at most quadratic growth under condition (ii) (see [9, Theorem 3.1]). However, they noted that their result cannot be applied to the equation
$\ddot{x}+\dot{x}+x^{3}=\dot{B}_{t}+\sin t,$ (1.4)
(see [9, Section 4, P. 342]). Their numerical simulations indicate that Eq. (1.4) does admit a periodic solution in distribution, leaving this case as an open problem. The authors further suggested that stochastic Levinson’s conjecture with more complex coefficients is an interesting problem for future research.
Since then, the stochastic Levinson conjecture has been widely studied. Jiang and Li [8] established sufficient conditions for the conjecture to hold, assuming the potential function exhibits at most quadratic growth and that the friction matrix is independent of $x,\dot{x}$. Ji et al. [10] investigated the existence of periodic probability solutions for stochastic differential equations (SDEs) with irregular coefficients using time-periodic Fokker-Planck equations. Subsequently in [11], they demonstrated the convergence of these solutions for SDEs with less regular coefficients. As an application, both papers [10, 11] provided sufficient conditions for the stochastic Levinson conjecture to hold, leveraging the Lyapunov function constructed by Wu [22]. For the theories of more complex time-recurrence solutions in distribution, we can refer to [3, 19, 4, 7, 18] and their many references therein.
Notably, system (1.3) encompasses the time-periodic Langevin equation
$\ddot{x}+\gamma(t)\,\dot{x}+\nabla_{x}V(x,t)=\dot{B}_{t},$ (1.5)
and time-periodic stochastic damping Hamiltonian systems (see [22]). Feng, Zhao, and Zhong [5, Theorem 4.17] further proved that the Langevin equation with constant positive friction admits a unique periodic solution in distribution.
This paper aims to establish the existence of periodic solutions in distribution for stochastic time-periodic Newtonian systems with either relaxation oscillating friction, or uniformly positive friction, or beyond.
We first present a general existence criterion for distributed periodic solutions to stochastic time-periodic Hessian-driven friction Newtonian systems in which either the potential force or the friction exhibits at least polynomial growth at infinity:
$\begin{cases}dx_{t}=y_{t}dt,\\ dy_{t}=-\big{[}D^{2}F(x_{t})\,y_{t}+\nabla_{x}V(x_{t},t)+E(x_{t},y_{t},t)\big{]}\,dt+\Sigma(x_{t},y_{t},t)\,d{B}_{t},\end{cases}$ (1.6)
which is stated as follows.
Theorem 1.2
The system (1.6) admits at least a $\mathcal{T}$-periodic solution in distribution if it satisfies Hypotheses (H1)-(H5). Here
(H1) All coefficients in system (1.6) are continuous in their respective domains and locally Lipschitz continuous in $(x,y)$. Let $F\in C^{2}(\mathbb{R}^{n}),V(x,t)\in C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n}\times\mathbb{R}^{+})$, and $E(x,y,t)$, $\Sigma(x,y,t)$ also be periodic in time $t$ with period $\mathcal{T}$. Additionally, $V(x,t)$ is lower bounded.
(H2) There exists a constant $a>0$ such that
$\lim_{\|x\|\rightarrow+\infty}\Big{[}V(x,t)+aF(x)-\frac{a^{2}}{2}\|x\|^{2}\Big{]}=+\infty,\qquad\hbox{uniformly in}~{}t\in\mathbb{R}^{+}.$
(H3) There exist positive constants $b,M$ and $m$ such that the friction-potential gradient inner product satisfies
$-\langle\nabla_{x}V(x,t),\nabla F(x)-ax\rangle\leq M-b\|x\|^{2m},\qquad~{}\forall(x,t)\in\mathbb{R}^{n}\times\mathbb{R}^{+}.$
(H4) $E(x,y,t)$ is bounded by $e$ and
$\lim_{\|x\|\rightarrow+\infty}\frac{e\|\nabla F(x)-ax\|}{\|x\|^{2m}}=0.$ (1.7)
(H5) There exist positive constants $c_{1},c_{2},M_{1}$ and $M_{2}$ with $(c_{1}+c_{2})<1$ such that
$|V_{t}(x,t)|\leq c_{1}\,b\|x\|^{2m}+M_{1},\qquad\quad\qquad\qquad\qquad~{}~{}\forall(x,t)\in\mathbb{R}^{n}\times\mathbb{R}^{+},$
$\|\Sigma\|_{2}^{2}:={\rm Tr}[(\Sigma\Sigma^{\top})(x,y,t)]\leq 2c_{2}\,(a\|y\|^{2}+b\|x\|^{2m})+M_{2},\quad\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+}.$
Let $p$ and $q$ be integers and consider the simplest case where $n=1$,
$V(x)=a_{2p}x^{2p}+a_{2p-1}x^{2p-1}+\cdots+a_{1}x+a_{0},\ F(x)=c_{2(q+1)}x^{2(q+1)}+c_{2q+1}x^{2q+1}+\cdots+c_{1}x+c_{0},$
$E(x,y,t)\equiv e(t)\equiv e(t+\mathcal{T})$ and $\Sigma\equiv 0$. The necessary condition for system (1.1) to be dissipative or the existence of a $\mathcal{T}$-periodic solution is
$a_{2p}>0,\qquad c_{2(q+1)}>0.$ (1.8)
Theorem 1.1 confirms such a condition is also sufficient when $p>q$.
To demonstrate the strength of Theorem 1.2, we examine how far it can be applied to system (1.6) with
$F(x)=Q_{2q}(x)+\sum_{k=2}^{2q-1}Q_{k}(x),\qquad V(x)=P_{2p}(x)+\sum_{j=1}^{2p-1}P_{j}(x),$
where each $Q_{k}(x)$ and $P_{j}(x)$ is a homogeneous polynomial of degree $k$ and $j$, respectively. In higher dimensions, the scalar dissipative condition (1.8) corresponds to the requirement that
the leading terms $P_{2p}(x)$ and $Q_{2q}(x)$ are positive definite.
Under this minimal assumption alone, we use Theorem 1.2 to prove the stochastic Levinson conjecture holds, which greatly improves upon the counterpart result in [15]:
Theorem 1.3
Let $E(x,y,t),\Sigma(x,y,t)$ be continuous, $\mathcal{T}$-periodic in time $t$, locally Lipschitz continuous with respect to $(x,y)$, and $E(x,y,t)$ be bounded. Assume that
$\|\Sigma\|_{2}^{2}(x,y,t)\leq c\big{(}\|y\|^{2}+\|x\|^{2(p+q-1)}\big{)}+C,$ (1.9)
where $c$ is given in (3.10). If both $Q_{2q}(x)$ and $P_{2p}(x)$ are positive definite and either $Q_{2q}(x)=\|x\|^{2q}$ or $P_{2p}(x)=\|x\|^{2p}$, then system (1.6) admits a $\mathcal{T}$-periodic solution in distribution.
Note that the time-periodic Langevin equation (1.5) is included in general equation:
$\begin{cases}dx_{t}=y_{t}dt,\\ dy_{t}=-\left[C(x_{t},y_{t},t)y_{t}+\nabla V(x_{t})+E(x_{t},y_{t},t)\right]\,dt+\Sigma(x_{t},y_{t},t)\,d{B}_{t}.\end{cases}$ (1.10)
Finally, we prove the existence of a periodic solution in distribution for (1.10) with a uniformly positive friction and a potential of degree exceeding $2$.
Theorem 1.4
The system (1.10) admits at least a $\mathcal{T}$-periodic solution in distribution if it satisfies the Assumptions (A1)-(A4). Here
(A1) All coefficients in system (1.10) are continuous in their respective domains and locally Lipschitz continuous in $(x,y)$. Let $V(x)\in C^{2}(\mathbb{R}^{n})$ be lower bounded and $C(x,y,t),E(x,y,t)$ $\Sigma(x,y,t)$ also be periodic with respect to time $t$ with period $\mathcal{T}$.
(A2) There exist positive constants $\alpha$ and $\beta$ such that
$C^{s}(x,y,t)\geq 2\alpha I_{n};\ \ \|C(x,y,t)\|\,\vee\,|E(x,y,t)\|\leq\beta,\qquad\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+},$
where $C^{s}(x,y,t):=\frac{1}{2}\left(C(x,y,t)+C^{\top}(x,y,t)\right)$.
(A3) There exist positive constants $b,M$ and $\epsilon$ such that
$\langle x,\nabla V(x)\rangle\geq b\|x\|^{2+\epsilon}-M,\qquad\quad\forall x\in\mathbb{R}^{n}.$
(A4) There exist positive constants $c$ and $M_{1}$ with $c<1$ such that
$\|\Sigma\|_{2}^{2}(x,y,t)\leq c\alpha\,(\|y\|^{2}+b\|x\|^{2+\epsilon})+M_{1},\qquad\qquad~{}\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+}.$
The proofs of Theorems 1.2-1.4 rely on Khasminskii’s criterion [12, Theorem 3.8] via a delicate Lyapunov function construction, building upon ideas from Jiang et al  [^1] On the exponential ergodicity and large deviations of stochastic Hessian-driven damping Hamiltonian systems. Preprint.. Our polynomial Lyapunov functions are smaller than the exponential functions introduced by Wu [22]. Polynomial functions can offer superior probabilistic properties: their slower growth reduces the possibility of exponential explosion and broadens the admissible class of initial distributions.
A key contribution of this paper is the Lyapunov function defined in (3.1), which combines mixed energy, potential energy, and a friction term. This formulation enables considerably more concise computations and provides the following advantages:
  • Stochastic Levinson’s Conjecture Holds Under Broad Conditions
    All known results establishing the stochastic Levinson conjecture require the friction matrix to be uniformly positive definite, the potential force to diverge to infinity and the noise to be bounded. Moreover, [9] and [8] additionally impose the assumption that the potential function $V(x)$ exhibits at most quadratic growth under condition (ii) of Theorem 1.1. The result in [5] specifically addresses the case where the friction matrix is a positive constant multiple of the identity matrix. Works such as [10, 11] accommodate a broader class of friction matrices that are unbounded in $x$ and bounded in $y$, but extracting the unbounded part $\Phi(x)$ of the friction matrix $C(x,y,t)$ required for their third assumption remains challenging.
    In contrast, our main Theorem 1.2 proves that distributed periodic solutions exist if $V(x,t)+aF(x)\to\infty$ uniformly in $t$ as $\|x\|\to\infty$ and the friction-potential gradient inner product grows at least like an even-power polynomial. The result holds even under bounded friction matrix perturbations (requiring higher potential growth). This result confirms the stochastic Levinson conjecture under broader conditions, including cases not covered previously:
    1. 1.
      stochastic relaxation oscillator: alternating friction definiteness;
    2. 2.
      the Newton systems with friction matrix no lower bound;
    3. 3.
      unrestricted growth rate in the friction-potential relationship;
    4. 4.
      the Newton systems with bounded potential force;
    5. 5.
      unbounded noises, potentially;
    6. 6.
      bounded perturbations of Hessian friction: avoiding unbounded extraction.
  • Distributed Periodic Solutions Always Exist for Periodically Forced and Noisy Polynomial Systems
    Theorem 1.3 establishes the existence of distributed periodic solutions for periodically forced and noise-driven polynomial systems with positive definite leading terms, which is a minimal assumption because it is the necessary condition for dissipation.
    Theorem 1.1 requires strong dissipativity: the second inequality in (1.2) implies that the potential grows at least quadratically at infinity and has a higher degree than the friction (i.e., $p>q$). Consequently, it applies only to strongly dissipative deterministic time-periodic Newtonian systems. This excludes seminal cases like the time-periodically forced van der Pol equation — a milestone in dynamical systems (see [25, 26, 27]). Furthermore, the first inequality in (1.2) mandates that friction admits a positive lower bound.
    By contrast, Theorem 1.3 applies to both deterministic and stochastic time-periodic Newtonian systems, with no restrictions on the relative degrees of friction and potential. This broader applicability significantly extends Theorem 1.1, even in the scalar case (see [15, P. 2619-2620]). Specifically, it covers periodically forced and noise-driven van der Pol oscillator, van der Pol-Duffing oscillator, and their higher-dimensional counterparts. It even applies to dissipative systems with friction lacking a lower bound.
    Moreover, Theorem 1.3 resolves Open Problem (1.4) posed in [9, Section 4].
  • Relaxing Infinity Requirement of Potential in Plasma Physics
    Luigi Nocera (MR4013829)(see his review on [10]) observed that existing results require the potential force to be infinite at infinity, which excludes certain plasma physics models. He suggested that distributed periodic solutions should exist for these forces and that the requirement could likely be relaxed for relevant plasma models.
    Theorem 1.2 achieves precisely this relaxation: even when $\nabla_{x}V(x,t)$ is bounded, condition (H3) holds provided $\nabla F(x)$ is so large that their gradient inner product grows like still an even-power polynomial, as demonstrated in (4.1)-(4.3) of Examples 13. This further demonstrates the power of our Lyapunov function approach.
The organization of this paper is as follows. In Section 2, necessary preliminaries are given. In Section 3, we will give the proofs of Theorem 1.2-1.4. In Section 4, we will provide a series of applications originating from plasma physics. In Section 5, we summarize our main results and state our further perspectives on stochastic time-recurrent Newtonian systems.
2 Preliminaries
This section is devoted to provide preliminaries for our setup. Firstly, the definition of periodic solution in distribution is introduced, which is referred to [3, 8, 9, 12, 10, 11, 5, 16, 17]. Then, we present Khasminskii’s criterion in [12, Theorem 3.8], which provides simple conditions for the existence of periodic solutions of SDEs in terms of Lyapunov functions, which is generalized by Ji et al. [10, 11] .
At first, we introduce the following notation. Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\in\mathbb{R}^{+}},\mathbb{P})$ be a filtered probability space satisfying the usual conditions. For periodic stochastic systems, the period of function $f$ with respect to time $t$ being $\mathcal{T}$ means that $f(x,t+\mathcal{T})=f(x,t)$ holds for all $t\in\mathbb{R}^{+}$ and $x\in\mathbb{R}^{n}$. Denote by $C_{\mathcal{T}}(\mathbb{R}^{n}\times\mathbb{R}^{+})$ the space of $\mathcal{T}$-periodic and continuous functions on $\mathbb{R}^{n}\times\mathbb{R}^{+}$ and by $C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n}\times\mathbb{R}^{+})$ the space $C^{2,1}(\mathbb{R}^{n}\times\mathbb{R}^{+})\cap C_{\mathcal{T}}(\mathbb{R}^{n}\times\mathbb{R}^{+})$.
Consider the following $\mathcal{T}$-periodic stochastic differential equation
$\begin{cases}dZ_{t}=b(Z_{t},t)\,dt+\sigma(Z_{t},t)\,dB_{t},~{}~{}\quad t>s,\\ Z_{s}=z,\end{cases}$ (2.1)
where the drift $b:\mathbb{R}^{n}\times\mathbb{R}^{+}\to\mathbb{R}^{n}$ and noise intensity $\sigma:\mathbb{R}^{n}\times\mathbb{R}^{+}\to\mathbb{R}^{n\times n}$ are $\mathcal{T}$-periodic in their second variable $t$. Let $Z^{s,z}(t)$ be the solution of (2.1) and ${\mathbb{P}}(s,z,t,A):=\mathbb{P}\big{(}Z^{s,z}(t)\in A\big{)}$ its transition probability function. Then ${\mathbb{P}}(s,z,t,A)$ is $\mathcal{T}$-periodic in the sense that ${\mathbb{P}}(s+\mathcal{T},z,t+\mathcal{T},\cdot)\equiv{\mathbb{P}}(s,z,t,\cdot)$ for all $z\in\mathbb{R}^{n}$ and $0\leq s\leq t$, see [12, Theorem 3.4]. We call a Markov process $Z(t)$ to be $\mathcal{T}$-periodic if its law ${\mathbb{P}}_{0}(t,A):={\mathbb{P}}\{Z(t)\in A\}$ satisfies the equation
${\mathbb{P}}_{0}(s,A)=\int_{\mathbb{R}^{n}}{\mathbb{P}}_{0}(s,dz)\,{\mathbb{P}}(s,z,s+\mathcal{T},A)\equiv{\mathbb{P}}_{0}(s+\mathcal{T},A),\quad\forall s\geq 0,\ A\in\mathcal{B}(\mathbb{R}^{n}),$ (2.2)
see [12, P. 64,  (3.17)]. (2.2) is equivalent to for any $s\geq 0$, the laws of $Z(s)$ and $Z(s+\mathcal{T})$ are identical. According to [8, 9], such a $\mathcal{T}$-periodic solution $Z(t)$ is called to be periodic solution in distribution or distributed periodic solution.
Now, we give the Khasminskii criterion in [12, Theorem 3.8], which implies the existence of $\mathcal{T}$-periodic solution of Eq. (2.1). Moreover, it plays a crucial role in the proof of Theorem 1.2.
Theorem 2.1
([12, Theorem 3.8]) Suppose that the coefficients of (2.1) are $\mathcal{T}$-periodic in time $t$ and satisfy the local Lipschitz conditions, and suppose further that there exists a non-negative function $\Psi(x,t)\in C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n},\mathbb{R}^{+})$, which satisfies
$\lim_{R\to\infty}~{}\inf_{\|x\|>R,~{}t\in\mathbb{R}^{+}}\Psi(x,t)=+\infty,$ (2.3)
and
$\lim_{R\to\infty}~{}\sup_{\|x\|>R,~{}t\in\mathbb{R}^{+}}\mathcal{L}\Psi(x,t)=-\infty,$ (2.4)
where the operator $\mathcal{L}$ is the generator of (2.1). Then (2.1) admits a $\mathcal{T}$-periodic Markov process.
It should be noted that there are several different ways of defining periodic solutions, but they are essentially the same. To the best of our knowledge, the earliest work on the periodic solution of stochastic processes can be traced back to Khasminskii [12, P. 64]. Ji et al. [10, 11] defined the stochastic periodic solution in terms of Fokker-Planck equations, while Feng et al. [5] defined it in terms of semigroups.
3 The Proof of Main Theorems
Proof of Theorem 1.2.  Define the Lyapunov function
$\Psi(x,y,t):=\frac{1}{2}\|y+\nabla F(x)-ax\|^{2}+\Big{[}V(x,t)+aF(x)-\frac{a^{2}}{2}\|x\|^{2}\Big{]}+D.$ (3.1)
Here, $a>0$ is given in Hypothesis (H2) and related to convergent rate, and $D$ is a sufficiently large constant such that $\Psi\geq 1$ for any $(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+}$. The first term is called to be the mixed energy. The selection of cross term between variables $x$ and $y$ is absolutely crucial when constructing Lyapunov functions for damped Newtonian systems. We choose the cross term to be $\langle y,\nabla F(x)-ax\rangle$ within the mixed energy, leading to significantly more concise computations and greatly advancing the applicable scope as described in Introduction.
Obviously, function $\Psi$ is $\mathcal{T}$-periodic and satisfies by Hypothesis (H2)
$\inf_{t\in\mathbb{R}^{+}}\Psi(x,y,t)\to+\infty,$
uniformly in time $t$ as $\|x\|+\|y\|$ tends to +$\infty$+ ∞. This implies function $\Psi$ satisfies condition (2.3) of Theorem 2.1.
The generator of system (1.6) is given by: for any $f\in C^{1,2,1}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+}\to\mathbb{R}$,
$\mathcal{L}f(x,y,t)$ (3.2)
$=$ $\partial_{t}f(x,y,t)+\frac{1}{2}\sum_{i,j=1}^{n}(\Sigma\Sigma^{\top})_{i,j}(x,y,t)\,\partial^{2}_{y_{i}\,y_{j}}f(x,y,t)$
$+\langle y,\nabla_{x}f(x,y,t)\rangle-\langle D^{2}F(x)\,y+\nabla_{x}V(x,t)+E(x,y,t),\nabla_{y}f(x,y,t)\rangle.$
It remains to check condition (2.4) of Theorem 2.1. By computation, we have
$\nabla_{y}\Psi=y+\nabla F(x)-ax,$ (3.3)
$\nabla_{x}\Psi$ $=$ $[\nabla_{x}V(x,t)+a\nabla F(x)-a^{2}x]$ (3.4)
$+[D^{2}F(x)y-ay+D^{2}F(x)\nabla F(x)-a\nabla F(x)-aD^{2}F(x)x+a^{2}x]$
$=$ $\nabla_{x}V(x,t)+D^{2}F(x)\,y-ay+D^{2}F(x)\nabla F(x)-aD^{2}F(x)\,x.$
Then using (3.3) and (3.4), we get
$\langle y,\nabla_{x}\Psi(x,y,t)\rangle-\langle D^{2}F(x)\,y+\nabla_{x}V(x,t),\nabla_{y}\Psi(x,y,t)\rangle$ (3.5)
$=$ $-\big{[}a\|y\|^{2}+\langle\nabla_{x}V(x,t),\nabla F(x)-ax\rangle\big{]}.$
Substituting (3.5) into (3.2) with $f=\Psi$, and then applying Hypotheses (H3)-(H5), we obtain that
$\mathcal{L}\Psi(x,y,t)$ (3.6)
$=$ $\partial_{t}V(x,t)+\frac{1}{2}\|\Sigma\|_{2}^{2}(x,y,t)-\langle E(x,y,t),y+\nabla F(x)-ax\rangle$
$-a\|y\|^{2}-\langle\nabla_{x}V(x,t),\nabla F(x)-ax\rangle$
$\leq$ $(c_{1}+c_{2}-1)\,(a\|y\|^{2}+b\|x\|^{2m})+\Big{(}M_{1}+\frac{1}{2}M_{2}+M\Big{)}$
$+e\|y\|+e\|\nabla F(x)-ax\|$
$\leq$ $\frac{1}{2}(c_{1}+c_{2}-1)\,(a\|y\|^{2}+b\|x\|^{2m})+\Big{(}M_{1}+\frac{1}{2}M_{2}+M+M^{*}\Big{)},$ (3.7)
for some positive constant $M^{*}$, where the first inequality (3.6) has used Hypotheses (H5), (H3) and $\|E(x,y,t)\|\leq e$, while the second inequality (3.7) has used Hypothesis (H4). This demonstrates (2.4) from the fact that $10<c_{1}+c_{2}<$ 1 and then the conclusion follows immediately from Theorem 2.1. $\Box$
Proof of Theorem 1.3.  The proof of this theorem is mainly obtained by applying Theorem 1.2, so we only need to verify Hypotheses (H1)-(H5) one by one. The proof is divided into several steps. It is easy to see that Hypothesis (H1) is satisfied by our assumption here.
Step 1. We first prove that Hypothesis (H2) holds. In order to check it, we have to find a positive constant $a$ such that
$\lim_{\|x\|\rightarrow+\infty}\Big{[}V(x)+aF(x)-\frac{a^{2}}{2}\|x\|^{2}\Big{]}=+\infty.$ (3.8)
Indeed, if either $p>1$ or $q>1$, then by the positive definiteness of $Q_{2q}(x)$ and $P_{2p}(x)$, (3.8) holds for any given positive $a$. It remains to consider the case $p=q=1$. In this case, $Q_{2}(x)$ and $P_{2}(x)$ are positive definite quadratic forms. Thus, there exists a positive constant $\lambda$ such that
$Q_{2}(x)\,\wedge\,P_{2}(x)\geq\lambda\|x\|^{2}.$
If necessary, we can shrink $\lambda$ so that
$\Big{[}V(x)+aF(x)-\frac{a^{2}}{2}\|x\|^{2}\Big{]}\geq\Big{(}\lambda+\lambda a-\frac{a^{2}}{2}\Big{)}\,\|x\|^{2},$
holds. Choose $a\in(0,2\lambda)$ so small that $(\lambda+\lambda a-\frac{a^{2}}{2})>0$. Such an $a$ guarantees that (3.8) holds. Therefore, there always exists a positive constant $a$ such that (3.8) holds, which proves that Hypothesis (H2) is always true.
Step 2. We then claim that there exist positive constants $b$ and $M$ such that
$\langle\nabla V(x),\nabla F(x)-ax\rangle\geq b\|x\|^{2(p+q-1)}-M,\qquad\forall x\in\mathbb{R}^{n}.$ (3.9)
This implies that Hypothesis (H3) holds.
Without loss of generality, we assume that $P_{2p}(x)=\|x\|^{2p}$. First, let $p=q=1$. Then by the Euler formula, we can obtain
$\langle\nabla V(x),\nabla F(x)-ax\rangle$
$=$ $\langle 2x+v_{0},\nabla Q_{2}(x)-ax\rangle$
$=$ $4Q_{2}(x)-2a\|x\|^{2}+\langle v_{0},\nabla Q_{2}(x)\rangle-a\langle v_{0},x\rangle$
$\geq$ $2(2\lambda-a)\|x\|^{2}+\langle v_{0},\nabla Q_{2}(x)\rangle-a\langle v_{0},x\rangle$
$\geq$ $(2\lambda-a)\|x\|^{2}-M,$
for some positive constant $M$, which implies (3.9) holds in this case.
In the remaining case $p+q>2$, it follows from the positive definiteness of $Q_{2q}(x)$ that there is a positive constant $\nu$ such that $Q_{2q}(x)\geq\nu\|x\|^{2q}$ for all $x\in\mathbb{R}^{n}$. Employing the Euler formula, we deduce that
$\langle x,\nabla Q_{2q}(x)\rangle=2q\,Q_{2q}(x)\geq 2q\nu\|x\|^{2q},$
and further
$\langle\nabla V(x),\nabla F(x)-ax\rangle$
$=$ $\big{\langle}2p\|x\|^{2(p-1)}x+\sum_{j=1}^{2p-1}\nabla P_{j}(x),\nabla Q_{2q}(x)+\sum_{k=2}^{2q-1}\nabla Q_{k}(x)-ax\big{\rangle}$
[…]
$\geq$ $4pq\,\nu\|x\|^{2(p+q-1)}+o(\|x\|^{2(p+q-1)})\ \quad({\rm as}\ \|x\|\rightarrow+\infty)$
$\geq$ $2pq\,\nu\|x\|^{2(p+q-1)}-M.$
This proves (3.9) for the case $p+q>2$.
Step 3. Finally, we verify Hypotheses (H4) and (H5). Note that a positive constant $c$ in (1.9) should satisfy
$c<\begin{cases}{\rm min}\{2(2\lambda-a),2a\},\ {\rm if}\ p=q=1,\\ {\rm min}\{4pq\nu,2a\},\ {\rm if}\ p+q>2.\end{cases}$ (3.10)
Thus far, we have verified Hypotheses (H1)-(H4). It is straightforward to verify that Hypothesis (H5) holds for $m=2(p+q-1)$.
Based on the above displays, the conclusion of this theorem follows immediately from Theorem 1.2. $\Box$
Remark 3.1
In the deterministic situation, Li et al [15] considered
$\ddot{x}+D^{2}F(x)\dot{x}+\nabla V(x)=e(t)\equiv e(t+\mathcal{T}),$ (3.11)
and gave sufficient conditions for system (3.11) to admit a $\mathcal{T}$-periodic solution. For scalar case, that is, $n=1$, and the friction and the potential are polynomials:
$V(x)=a_{0}+a_{1}x+a_{2}x^{2}+\cdots a_{2p}x^{2p}(a_{2p}>0),\ F(x)=c_{0}+c_{1}x+c_{2}x^{2}+\cdots c_{2(q+1)}x^{2(q+1)}(c_{2(q+1)}>0),$
they proved that system (3.11) has a $\mathcal{T}$-periodic solution if $p>q\geq 0$. For higher dimensional system, their result applies to $F(x)=\|x\|^{2q},\ V(x)=\|x\|^{2p}$ to conclude that the existence of a $\mathcal{T}$-periodic solution if $p>q>0$. However, it is worth noting that their result cannot be applied to the well-known periodically forced van der Pol equation and its higher-dimensional version.
Note that system (3.11) with $n=1$ is dissipative if and only if $a_{2p}>0$ and $c_{2(q+1)}>0$. For higher-dimensional systems, this condition corresponds to the leading terms of $V(x)$ and $F(x)$ being positive definite. Under this minimal assumption, we proved in Theorem 1.3 that both systems (3.11) and (1.6) admit a $\mathcal{T}$-periodic solution. This result significantly improves upon those in [15].
Proof of Theorem 1.4. Setting $F(x):=\alpha\|x\|^{2},a:=\alpha$ in (3.1) leads to the following Lyapunov function
$\Psi(x,y)=\frac{1}{2}\|y+\alpha x\|^{2}+V(x)+\frac{\alpha^{2}}{2}\|x\|^{2}.$
By Hypothesis (A1), it is easy to see that
$\lim_{\|x\|+\|y\|\rightarrow+\infty}\Psi(x,y)=+\infty.$
This implies the function $\Psi$ satisfies condition (2.3) of Theorem 2.1. In addition, using Hypotheses (A2)-(A4), we have
$\mathcal{L}\Psi=$ $-\alpha[\|y\|^{2}+\langle x,\nabla V\rangle(x)]-\langle y,(C^{s}-2\alpha I)y\rangle-\alpha\langle x,(C-2\alpha I)y\rangle-\langle E,y+\alpha x\rangle+\frac{1}{2}\|\Sigma\|_{2}^{2}$
$\leq$ $-\alpha\big{[}\|y\|^{2}+\langle x,\nabla V(x)\rangle\big{]}+\alpha(\beta+2\alpha)\|x\|\,\|y\|+\beta(\|y\|+\alpha\|x\|)+\frac{1}{2}\|\Sigma\|_{2}^{2}$
$\leq$ $-\alpha\big{[}\|y\|^{2}+\langle x,\nabla V(x)\rangle\big{]}+\frac{\alpha}{4}\|y\|^{2}+\alpha(\beta+2\alpha)^{2}\|x\|^{2}+\beta(\|y\|+\alpha\|x\|)+\frac{1}{2}\|\Sigma\|_{2}^{2}$
$\leq$ $-\frac{3\alpha}{4}\|y\|^{2}+\beta\|y\|-\alpha b\|x\|^{2+\epsilon}+\alpha(\beta+2\alpha)^{2}\|x\|^{2}+\alpha\beta\|x\|+\alpha M+\frac{1}{2}\|\Sigma\|_{2}^{2}$
$\leq$ $-\frac{\alpha}{2}\left(\|y\|^{2}+b\|x\|^{2+\epsilon}\right)+M^{*}+\frac{1}{2}\|\Sigma\|_{2}^{2}$
$\leq$ $-\frac{\alpha}{2}(1-c)\left(\|y\|^{2}+b\|x\|^{2+\epsilon}\right)+M^{*}+\frac{M_{1}}{2},$
which implies the function $\Psi$ satisfies condition (2.4). The conclusion follows immediately from Theorem 2.1. $\Box$
Remark 3.2
We note that Ji et al [10, 11] have established Theorem 1.4 under bounded noise conditions when the friction matrix is bounded, using the exponential-type Lyapunov function introduced in [22]. However, our approach employs a polynomial Lyapunov function. A polynomial function offers superior probabilistic properties compared to an exponential one. Specifically, its slower growth rate mitigates the risk of exponential explosion and expands the admissible space of initial distributions.
4 Plasma Physics Models with Bounded Potential Force
This section is motivated by Luigi Nocera’s remarks on [10], see MR4013829. He observed that existing results require potentials to be “infinite at infinity,” which excludes applications to certain plasma physics models. Nocera suggested that distributed periodic solutions should exist for these models and that the requirement could likely be relaxed. Our Theorem 1.2 achieves precisely this relaxation.
We present a series of examples with potential functions drawn from [21, 13, 20] (sometimes slightly modified). In these cases, the potential force is bounded, while friction and noise terms remain unbounded, which further highlights the strength of our theorem. Throughout, we assume the diffusion coefficient $\Sigma$ satisfies the standard assumptions in Hypothesis (H1).
Example 1
Suppose that
$V(x,t)=\ln(2+\sin t+\|x\|^{2}),\qquad F(x)=\|x\|^{4}+\|x\|^{2}.$
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{3}(\|x\|^{2}+\|y\|^{2})+C$ with $c_{3}<16$, then it admits a $2\pi$-periodic solution in distribution.
In fact, taking $a:=8$ and $D:=19$, we have
$\partial_{t}V(x,t)$ $=$ $\frac{\cos t}{2+\sin t+\|x\|^{2}},$
$\nabla_{x}V(x,t)$ $=$ $\frac{2x}{2+\sin t+\|x\|^{2}},$
$\nabla F(x)-8x$ $=$ $2x\,(2\|x\|^{2}-3).$
Obviously, $\|\nabla_{x}V(x,t)\|\leq 1$ and $D^{2}F(x)=(4\|x\|^{2}+2)I_{n}+8xx^{\top}$ is positive definite. Moreover,
$-\langle\nabla_{x}V(x,t),\nabla F(x)-8x\rangle=-\frac{4\|x\|^{2}\,(2\|x\|^{2}-3)}{2+\sin t+\|x\|^{2}}~{}\leq-8\|x\|^{2}+36.$ (4.1)
It is easy to see that Hypotheses (H1)-(H5) hold with $b=8,\,m=1,$ and $M=36$. $\Box$
Example 2
Assume that
$V(x,t)=\sqrt{2+\sin t+\|x\|^{2}},\qquad F(x)=\|x\|^{4}+\|x\|^{2}.$
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{4}(\|x\|^{3}+\|y\|^{2})+C$ with $c_{4}<4$, then it admits a $2\pi$-periodic solution in distribution.
Indeed, taking $a:=2$ and $D:=1$, we obtain
$\partial_{t}V(x,t)$ $=$ $\frac{\cos t}{2\sqrt{2+\sin t+\|x\|^{2}}},$
$\nabla_{x}V(x,t)$ $=$ $\frac{x}{\sqrt{2+\sin t+\|x\|^{2}}},$
$\nabla F(x)-2x$ $=$ $4\|x\|^{2}\,x.$
Meanwhile, $\|\nabla_{x}V(x,t)\|\leq 1$ and $D^{2}F(x)=(4\|x\|^{2}+2)I_{n}+8xx^{\top}$ is positive definite. In addition, it yields
$-\langle\nabla_{x}V(x,t),\nabla F(x)-2x\rangle=-\frac{4\|x\|^{4}}{\sqrt{2+\sin t+\|x\|^{2}}}\leq-2\|x\|^{3}+2.$ (4.2)
Hence, Hypotheses (H1)-(H5) hold with $b=2,\,m=3/2,$ and $M=2$. Then we conclude from Theorem 1.2 that there exists a $2\pi$-periodic solution in distribution. $\Box$
Example 3
Let
[…]
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{5}(\|x\|^{4}+\|y\|^{2})+C$ with $c_{5}<2$, then it admits a $2\pi$-periodic solution in distribution.
Indeed, taking $a:=1$ and $D:=2$, we have
$\partial_{t}V(x,t)$ $=$ $\cos t\,(1-\exp{(-\|x\|^{2})}),$
$\nabla_{x}V(x,t)$ $=$ $2(2+\sin t)\,\exp{(-\|x\|^{2})}\,x,$
$\nabla F(x)-x$ $=$ $\frac{1}{2}\exp{(\|x\|^{2})}\,(\|x\|^{2}+1)\,x-x.$
In addition, it follows that $\|\nabla_{x}V(x,t)\|\leq 3$ and
$D^{2}F(x)=\frac{1}{2}\exp(\|x\|^{2})\,(\|x\|^{2}+1)I_{n}+(\|x\|^{2}+2)\exp{(\|x\|^{2})}\,xx^{\top},$
is positive definite. Moreover, it yields
$-\langle\nabla_{x}V(x,t),\nabla F(x)-x\rangle$ (4.3)
$=$ $-(2+\sin t)\,\big{[}\|x\|^{4}+\|x\|^{2}-2\|x\|^{2}\,\exp{(-\|x\|^{2})}\big{]}$
$\leq$ $-\|x\|^{4}+3.$
Obviously, Hypotheses (H1)-(H5) hold with $b=1,\,m=2,$ and $M=3$. Then it yields from Theorem 1.2 that there exists a $2\pi$-periodic solution in distribution. $\Box$
Remark 4.1
[…]
5 Conclusions and Further Perspectives
This paper proves the existence of a distributed periodic solution for stochastic time-periodic Newton systems under the assumptions that both the friction and potential functions tend to positive infinity at infinity, and the friction-potential gradient inner product grows at least like an even-power polynomial. This result is robust regarding bounded perturbations of the Hessian-friction matrix, at the cost of a slight increase in the growth rate of the potential function at infinity. Consequently, our method can be applied to systems like the one in Eq. (1.6), where $V(x)=\|x\|^{6}$ and $D^{2}F(x)$ is replaced by $D^{2}\left(\|x\|^{4}\pm\|x\|^{2}\right)+B(x,y,t)$ with $B$ being bounded. To the best of our knowledge, none of the existing results can be applied to this system. Our results not only largely confirm the stochastic Levinson conjecture proposed in [9] and resolve the open problem and the interesting problem posed in [9, P. 342], but also possess advantages 1-6 enumerated in the Introduction, which are absent in existing results.
Our further research plan involves studying the uniqueness, exponential ergodicity, and exponential mixing of distributed time-recurrent solutions to the stochastic time-recurrent Newtonian equations, including distributed periodic solutions and quasi-periodic solutions, etc.
Acknowledgements  The project is supported by the National Natural Science Foundation of China with grant numbers 12171321.
Data Availability   Our manuscript has no associate data.
Declarations
Conflicts of Interest    All authors have no Conflict of interest.